Circular No. 7188

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

MICROLENSING EVENT TOWARDS THE LMC A. Becker, University of Washington, reports on behalf of the MACHO,

OGLE, and MPS collaborations on the progress of a high-magnification microlensing event, designated MACHO 99-LMC-2, regarding a star at α = 5^{h}16^{m}53^{s}.2, δ = −69^{o}16′30′′.2 (equinox 2000.0), which was constant in brightness (I = 19.9, R = 20.6, V = 21.0) since Jan. 1997 and was at a magnification of 16.3 (a brightening of 3 mag) as of 1999 June 4.4 UT. The best fit to microlensing indicates the object will reach a (currently unconstrained) maximum magnification near June 8.23 ± 0.6 day (a fit 2σ away provides a maximum magnification of 44.5 on June 7.94 ± 0.1 day). The brightening of the source allows for high-S/N photometry to constrain exotic microlensing phenomena (binary-source effects; ‘parallax’ effect), which lead to constraints on the lens location. High-S/N spectroscopy of the lensed source (presumably on the LMC main sequence) may become feasible, as it is expected to reach at least I = 15.8. This configuration also allows for enhanced planet-detection efficiency, if the lens harbors a planetary system. Observers should contact macho@astro.washington.edu or A. Becker at telephone 206-543-9205 to coordinate observations. A finding chart and data on this event may be found

at ftp://darkstar.astro.washington.edu/macho/Alert/99-LMC-2/.

V382 VELORUM B. Hidayat, M. Ikbal Arifyanto, J. Aria Utama, and S. Athiya, Boss-

cha Observatory, Lembang, write: “Spectroscopic observations of N Vel 1999 were obtained at the Bosscha Schmidt telescope (+ objective prism; dispersion 31.5 nm/mm at Hγ) on June 1.50 and 2.50 UT, with unfiltered IIa-O plates exposed for 12 and 6 min, respectively. Although the 12-min exposure yielded an overexposed spectrum, emission lines of H, He, and [O III] are clearly distinguishable. Based on the three spectra collected on June 2, the following lines are seen (positions and widths of the lines are based on a standard spectrum of a B9 star, secured with the same telescope): [O III] 500.7-nm (half the strength of Hβ), 495.9-nm (half the strength of Hβ), 436.3-nm (certainly observable); [Fe II] 492.4-nm (rather faint but traceable), 423.3-nm (faint); S II(?) 406.8-nm (rather broad, confuse, barely visible); [Ne III] 386.9-nm (half the strength of Hβ). All of the H lines down to H_{7} appear on our plates; their estimated width is ∼ 20.0 nm, decreasing rather slowly toward H_{7}. Shallow depression on the blue side of Hβ is noticable. He II 468.6-nm is rather wide.”

1999 June 4  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green